Q3. A marketer wants to market a product. The unit variable cost for producing this product is $16. The fixed cost is $400,000. The marketer expects to sell 80,000 units of the product and wants to earn a 30 percent markup on sales. How much should the markup price be for this product? (1 point, word limit: 50 words)
Answer #1

Let the sales be X . Markup on sales =X *.30 = .30X

Total Cost per unit =Variable cost per unit + Fixed cost per unit

                          = 16 + (400000/ 80000)

                         = 16 + 5

                         =$ 21 per unit

Markup = selling price -variable cost per unit

.30 X =X - 21

X -.30 X = 21

.70X =21

X = 21 /.70

     =$ 30 Per unit

markup price = $30 per unit